# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ IV.3 ESR in the FLT-driven regime
IV.4 Behavior of $\Omega$ and $T_{2}$ under DC bias
@@
In the following, we explore the behavior of the Rabi frequency $\Omega$ and coherence time $T_{2}$ as functions of the easily adjustable experimental parameter $eV_{\mathrm{DC}}$ for the out-of-plane situation studied in Figs. 3 and 5. The result for $T_{2}$ is shown in Fig. 7a while the one for the Rabi rates is depicted in Fig. 7b. We separate the contributions from the STT and FLT driven regimes in Fig. 7b, as these evolve quite differently under varying DC bias. The behavior of $T_{2}$ is relatively straightforward to understand based on the previous discussion. In the low-bias CB regime, the current through the impurity is not a full one-electron process. Since scattering between the transport electrons and the impurity is the primary source of decoherence in this model, the coherence time is maximized. Therefore, the CB regime effectively protects the impurity spin from the decoherence induced by the polarized current. As we approach the charging thresholds $\varepsilon$ and $\varepsilon+U$ getting out of the CB regime, $T_{2}$ rapidly decreases. Due to the single-orbital nature of our model, $T_{2}$ then reaches a constant and very small value once the DC bias $V_{\mathrm{DC}}$ exceeds these thresholds. This demonstrates that the choice of $V_{\mathrm{DC}}$ is critical for achieving favorable conditions for quantum control.
Fig. 7b illustrates the behavior of the Rabi rate in both the STT and FLT driven regimes. In the CB regime, the FLT Rabi rate dominates over the SST Rabi rates, following a logarithmic behavior, as indicated by the imaginary part of Eq. 25. For low biases, $\Omega_{\mathrm{FLT}}$ exhibits a linear dependence that does not vanish at zero DC bias, which explains the presence of an ESR signal at $V_{\mathrm{DC}}=0$ meV, even though the signal intensity is low due to the minimal current passing through the impurity. As the bias approaches the CB regime threshold, $\Omega_{\mathrm{FLT}}$ becomes highly nonlinear, reaching a maximum (minimum) at $\varepsilon+U$ ($\varepsilon$). According to Eq. 25, $\Omega_{\mathrm{FLT}}$ would theoretically diverge at $\varepsilon+U$ and $\varepsilon$ for $\hbar/\tau_{c}\ll\varepsilon+U,|\varepsilon|$, but finite temperature acts as a natural cutoff, preventing this divergence. Above the energy thresholds, the SST Rabi rate emerges strongly, leading to an incoherent ESR signal, as shown in Fig. 3 and 4. The sign of the Rabi rates in each driven regime determines the rotation direction: a negative sign induces clockwise rotation in the Bloch sphere, while a positive sign results in counterclockwise rotation.
@@
Figure 7c shows the figure of merit $\Omega T_{2}$, indicating a maximum of roughly 6 Rabi cycles before the steady state is reached for the FLT Rabi rate, as clearly depicted in Fig. 5. This underscores the importance of optimizing the DC bias to operate in the CB regime for achieving the best possible conditions for quantum control. Contrary, the $\Omega_{\mathrm{STT}}$ behavior only emerges outside the CB regime when the spin accumulation dominates the dynamics as Fig. 7b shows. Therefore, the product $\Omega_{\mathrm{STT}}T_{2}$ does not have a large value for any DC bias, as we inferred from Eq. 28. In the STT-driven regime, the spin is weakly driven, making this regime useless in terms of quantum control, but it can be utilized to polarize the QI to a high degree as we already showed in Fig. 3.
IV.5 Full DC voltage mapping and energy dressing
@@
In this section, we examine the ESR signal response under varying DC bias for a fixed polar angle of $\theta=\pi/18$. The results are presented in Fig. 8, where a 2D color map highlights the normalized ESR signal relative to the background (baseline off-resonance) current, preserving the characteristic peak and dip behaviors. The efficiency of the ESR signal closely follows the figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ shown in Fig. 7b, demonstrating low efficiency outside the CB regime beyond the energy thresholds $\varepsilon$ and $\varepsilon+U$. As the DC bias approaches zero, the background current diminishes, albeit at a different rate compared to $\Delta I$. This results in a non-zero ESR signal at zero bias, a phenomenon reported previously [19, 30]. This observation aligns with the earlier discussion that $\Omega_{\mathrm{FLT}}$ does not vanish at zero bias but rather at the electron-hole symmetry point, $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$. Indeed, there the signal completely disappears, as predicted by the exchange field formula, Eq. 23, and observed in Fig. 7.
Figure 8: Shifting of the ESR resonance position with DC bias. The ESR signal intensity, normalized to the background current, is shown as a color map, revealing peaks at positive DC biases and dips at negative DC biases. As the voltage approaches the thresholds $\varepsilon$ and $\varepsilon+U$, the ESR efficiency decreases substantially. The observed frequency shift agrees with the prediction of Eq. LABEL:energy_dress and deviates by up to approximately 8% (about 1.3 GHz) from the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz.
@@
An intriguing phenomenon observed in Fig. 8 is the shifting of the ESR resonance position. We emphasize that in these simulations, the tip-sample distance is kept constant, therefore, the couplings remain unchanged while sweeping the DC bias. This dressing of the resonance energy arises from the projection of the exchange field onto the quantization axis of the spin and only vanishes at the electron-hole symmetry point or when the spin is initialized exactly perpendicular to the exchange field. In these cases, the resonance occurs at the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz. This frequency shift has been consistently reported in the literature  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] , with various explanations proposed. Recently, it was suggested to be a result of a change of the $g$-factor by the electric field of the DC bias [19]. However, the net effect of the DC bias on the magnetic moment found in experiments [25], supported by DFT simulations, is orders of magnitude too small and rules out $g$-factor variation as the leading cause. There are also considerations regarding Ref. [22], which rules out the $g$-factor as a contributing factor. A $g$-modulation would discard the utility of the Fe atoms to drive the single-spin 1/2 Ti atom, contradicting experimental evidence, a point we discuss further in Ref. [73]. This leaves the exchange field as the most plausible mechanism for this phenomenon, though minor contributions from other sources cannot be entirely ruled out.
@@
The observable shift in the resonance frequency was used as a direct experimental method to determine the impurity parameters $\varepsilon$ and $\varepsilon+U$ [25], aligning well with DFT simulations [72]. Moreover, the shift depicted in Fig. 8 is clearly proportional to $\Omega_{\mathrm{FLT}}$ shown in Fig. 7a. Thus, measuring the resonance shift offers a practical way to obtain the Rabi rate. This prediction could be validated using alternative Rabi rate estimation methods, such as the one presented in [23].
@@
IV.6 Vector field dependence of the Rabi rate and ESR signal
@@
The polar angle $\theta$ formed between the external magnetic field and tip spin polarization appears naturally in all quantities of interest and is closely related to measure an efficient ESR signal, as we observed by comparing Figs. 5c and 6c. As such, it is important to understand the behavior of our model with respect to the polar angle. Furthermore, since the static magnetic field angle is a parameter easily adjusted in some experiments, confirming the agreement we have between model and experiment will provide validity of the driving field emerging from the modulation of the exchange magnetic field and the hopping.
@@
Figure 9: Example of the ESR signal behavior on resonance, calculating the total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, at negative (dotted line) and positive (solid line) DC bias as a function of the polar angle $\theta$ formed between the magnetic field and the tip spin polarization. The maximum ESR signal occurs at a predominantly out-of-plane angle, as a result of homodyne detection. A sizable signal for an in-plane angle arises due to the spin accumulation being noticeable for these DC biases.
@@
As an example of the angle behavior we have Fig. 9, where we show the total ESR amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$ for the optimal DC bias values that maximizes the quality factor $\Omega_{\mathrm{FLT}}T_{2}$, see Fig. 7c. $\Delta I_{s}$ denotes the symmetric part of the ESR signal, while $\Delta I_{a}$ provides the asymmetric one. Both are extracted from fitting the raw ESR signal to a symmetric and an antisymmetric Lorentzian center on the resonance frequency (which includes the frequency shift or energy dressing).
At $\theta=0,\pi$, the ESR signal completely vanishes as the driving field and the spin are parallel. This is not found experimentally due to the tip likely forming an angle $\theta^{\prime}$ with the plane containing the spin. In other words, in general, the tip spin polarization does not lie exactly on $z$ in the experiments, as it does in the model. This can be accounted for in the simulations by tilting the plane in which the magnetic field lies by an angle $\theta^{\prime}$, relative to the spin polarization in $z$.
@@
Figure 10: Full DC and vector field dependence. a) 2D color map of the ESR total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, normalized to the background current keeping its sign, as a function of the external field angle and DC bias, highlighting regions of high signal efficiency. $\Delta I_{s}$ gives the symmetric part of the ESR signal, while $\Delta I_{a}$ provides the asymmetric one. At $\theta=0,\pi$ and along the electron-hole symmetry lines, no ESR signal is observed. The white horizontal dashed lines correspond to the plots shown in Fig. 9. b) Figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ as a function of the polar angle $\theta$ and DC bias, showing significant variation within the Coulomb blockade (CB) regime, proportional to $\sin\theta$, and a rapid drop-off outside of it. The red dotted line indicates the electron-hole symmetry line at $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$.
@@
Outside of those special polar angles, both negative and positive DC bias in Fig. 9 show a similar behavior for the ESR amplitude that follows roughly the homodyne detection, as the largest ESR amplitude is found predominately out of plane ($\theta\sim 10$ degrees), which matches experimental evidence [34, 21]. However, at 90 degrees, we encounter a minimum of the ESR amplitude which goes against the conventional ESR-STM measurement. Here, the homodyne detection tells us that the amplitude should be zero for in-plane angle as the ESR signal is proportional to the scalar product of the spin and the tip direction. As we already mentioned, we identify this discrepancy as a limitation of the model at capturing the ESR signal in the FLT-driven regime, which includes a small spin accumulation contribution towards the $z$ axis, providing a non-zero ESR signal. A more robust theory that fully integrates the sequential tunneling and cotunneling transport is needed to obtain the full picture, the former being responsible for creating the exchange field. It is worth noting that, since we are plotting the ESR amplitude $\Delta I_{p}$, we do not visualize the flipping of the signal happening when the angle is sweeping from an out-of-plane value to an in-plane one, going through a pure fano-profile ESR signal.
@@
Turning our sight to Fig. 10a, we observe the ESR amplitude dependence versus both the DC bias and the polar angle. The behavior with the voltage is similar to the one presented in Fig. 8 and independent of the angle, but without visualizing the resonance shifting since we are plotting the ESR amplitude on resonance. Instead, if we focus on the angle dependence, one can see again that the ESR signal will vanish in the case of parallel ($\theta=0^{\circ}$) and anti-parallel ($\theta=180^{\circ}$) orientations. As we pointed out in Fig. 9, the ESR amplitude has a local minimum for a perpendicular alignment ($\theta=\pm 90^{\circ}$) for any DC bias due to the spin accumulation, however, the lower the absolute value of the voltage is, the smaller the ESR amplitude is at the perpendicular alignment. This is because the spin accumulation is reduced even further the lower the DC bias is, see Fig. 2, even though the quality factor $\Omega_{\mathrm{FLT}}T_{2}$ maximizes at $\theta=90$ degrees as Fig. 10b indicates.
Additionally, Fig. 10b clearly shows the sinusoidal dependence of the Rabi rate on the polar angle, as experimentally reported in [21], since $T_{2}$ remains independent of the magnetic field orientation. The electron-hole symmetry line, $eV_{\mathrm{DC}}=\varepsilon+U/2$, is represented by a red dotted line in Fig. 10b, where both the Rabi rate and ESR signal vanish, consistent with Fig. 10a at $eV_{\mathrm{DC}}/|\varepsilon|=0.25$.
@@
Outside the Coulomb blockade (CB) regime, the quality factor drops to extremely low values, following the decrease in $T_{2}$ observed in Fig. 7a. Near zero DC bias, the ESR amplitude relative to the background exhibits a local maximum or minimum, depending on whether $\theta\in(-90^{\circ},90^{\circ})$ or falls within the range $[90^{\circ},270^{\circ}]$, respectively. This ESR signal behavior aligns with the fact that the Rabi rate remains nonzero at this bias, as discussed earlier.

### v2

@@ IV.3 ESR in the FLT-driven regime
IV.4 On the measurability of QI spin dynamics via the current
We wish to clarify the physical interpretation of the charge current as a probe of the quantum impurity (QI) spin state, particularly in the context of coherent dynamics. While the current results from tunneling processes that inevitably affect the occupancy of the QI, it nonetheless acts as a form of weak measurement of the QI spin. Specifically, in the CB regime, the impurity remains singly occupied for long durations, and tunneling events are rare and weakly perturbative. In this regime, the spin-dependent tunneling rates provide a continuous readout of the QI spin projection along the axis defined by the spin polarization of the lead. The average current is thus sensitive to the expectation value of the spin component parallel to this axis, a principle that underlies the interpretation of ESR-STM magneto-resistant contrast and resonance features in numerous experimental works [1, 5, 6, 23, 25]. Since in the CB regime, where the QI remains mostly singly occupied, tunneling-induced decoherence is rare and minimal, allowing for coherent spin evolution over timescales much longer than the average time between tunneling events.
This interpretation is analogous to weak projective measurements in quantum computing, where the measurement of a qubit’s state is performed via its coupling to a pointer observable, such as a dispersive resonator or a charge detector [80]. There, as in our case, the observable read out (current or voltage) does not project onto a pure eigenstate in a single shot, but rather accumulates statistically meaningful information over many repetitions or over long-time averages. Our theoretical simulations incorporate this principle: we compute the time-dependent current arising from the QI spin dynamics and extract measurable features (resonance position, linewidth, contrast) by ensemble averaging encoded in the populations, coherences or expectation values of the spin components.
The situation is essentially similar in the STT regime, but transport is now highly incoherent. Frequent tunneling events rapidly degrade spin coherence, as reflected in the short $T_{2}$ times obtained in our simulations. Although $T_{2}$ can be partially enhanced, the incoherent nature of STT-driven ESR persists. Nonetheless, the current remains sensitive to the quantum evolution of the spin, particularly under resonant driving. We therefore interpret the current not as a fully projective measurement of the spin state at each instant, but as a time-resolved, spin-selective probe whose averaged behavior reflects the underlying quantum dynamics. This perspective is fully consistent with experimental ESR-STM protocols  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]and provides a natural basis for using the current as both a readout and a feedback channel for coherent and incoherent spin manipulation.
Let us now take a closer look at the experimental measurements and what setup enables precise spin readout. For concreteness, we refer to Ref. [11]. In STM, Rabi oscillations and other coherent control experiments are implemented using pulsed-ESR, where the spin state is driven by nanoscale RF pulses at the resonance frequency. The readout is based on lock-in detection of the tunneling current, which alternates between two measurement cycles. In cycle A, the RF drive is applied for a finite duration before the spin is allowed to relax, while in cycle B only the DC current is measured, RF is off. Subtracting the signals from these two cycles isolates the contribution of the driven spin dynamics. Rabi oscillations are then reconstructed by systematically varying the RF pulse length in cycle A. Longer pulses rotate the spin further on the Bloch sphere before relaxation, producing a distinct modulation of the tunneling current. Each data point corresponds to the average over hundreds of repeated A–B cycles and sweeping the RF pulse length in the cycle A, reconstructs the full Rabi oscillation.
Homodyne detection ensures that the measured current reflects the spin component parallel to the tip’s magnetization axis. In our model, this detection operates only in the FLT regime. In contrast, STT relies on a different detection mechanism but it also encodes the spin state under an incoherent evolution [26]. In all cases, the extremely fast Larmor oscillations are not resolved due to the limited temporal resolution of the instruments. Consequently, FLT measurements capture only the slower Rabi oscillations, as shown in Figs. 5b and 6b, whereas STT measurements exhibit an exponential decay, i.e., without the fast precession seen in Figs. 3b and 4b. These results are in agreement with Ref. [26].
IV.5 Behavior of $\Omega$ and $T_{2}$ under DC bias
@@
In the following, we explore the behavior of the Rabi frequency $\Omega$ and coherence time $T_{2}$ as functions of the easily adjustable experimental parameter $eV_{\mathrm{DC}}$ for the out-of-plane situation studied in Figs. 3 and 5. The result for $T_{2}$ is shown in Fig. 7a while the one for the Rabi rates is depicted in Fig. 7b. We separate the contributions from the STT and FLT driven regimes in Fig. 7b, as these evolve quite differently under varying DC bias. The behavior of $T_{2}$ is relatively straightforward to understand based on the previous discussion. In the low-bias CB regime, the current through the impurity is not a full one-electron process. Since scattering between the transport electrons and the impurity is the primary source of decoherence in this model, the coherence time is maximized. Therefore, the CB regime effectively protects the impurity spin from the decoherence induced by the polarized current. As we approach the charging thresholds $\varepsilon$ and $\varepsilon+U$ exiting the CB regime, $T_{2}$ rapidly decreases. Due to the single orbital nature of our model, $T_{2}$ then reaches a constant and very small value once the DC bias $V_{\mathrm{DC}}$ exceeds these thresholds. This demonstrates that the choice of $V_{\mathrm{DC}}$ is critical to achieve favorable conditions for quantum control.
Fig. 7b illustrates the behavior of the Rabi rate in both the STT and FLT driven regimes. In the CB regime, the FLT Rabi rate dominates over the STT Rabi rates, following a logarithmic behavior, as indicated by the imaginary part of Eq. (25). For low biases, $\Omega_{\mathrm{FLT}}$ exhibits a linear dependence that does not vanish at zero DC bias, which explains the presence of an ESR signal at $V_{\mathrm{DC}}=0$ meV, even though the signal intensity is low due to the minimal current passing through the impurity. As the bias approaches the CB regime threshold, $\Omega_{\mathrm{FLT}}$ becomes highly nonlinear, reaching a maximum (minimum) at $\varepsilon+U$ ($\varepsilon$). According to Eq. (25), $\Omega_{\mathrm{FLT}}$ would theoretically diverge at $\varepsilon+U$ and $\varepsilon$ for $\hbar/\tau_{c}\ll\varepsilon+U,|\varepsilon|$, but finite temperature will act as a natural cut-off if considered, preventing this divergence. Above the energy thresholds, the STT Rabi rate emerges strongly, leading to an incoherent ESR signal, as shown in Fig. 3 and 4. The sign of the Rabi rates in each driven regime determines the rotation direction: a negative sign induces clockwise rotation in the Bloch sphere, while a positive sign results in counterclockwise rotation.
@@
Figure 7c shows the figure of merit $\Omega T_{2}$, indicating a maximum of roughly 6 Rabi cycles before the steady state is reached for the FLT Rabi rate, as clearly depicted in Fig. 5. This underscores the importance of optimizing the DC bias to operate in the CB regime to achieve the best possible conditions for quantum control. Contrary, the $\Omega_{\mathrm{STT}}$ behavior only emerges outside the CB regime when the spin accumulation dominates the dynamics, as Fig. 7b shows. Therefore, the product $\Omega_{\mathrm{STT}}T_{2}$ does not have a large value for any DC bias, as we inferred from Eq. (28). Therefore, the STT-driven regime is inadequate for quantum control, but it can be utilized to polarize the QI to initialize the spin, as we already showed in Fig. 3.
IV.6 Full DC voltage mapping and energy dressing
@@
In this section, we examine the ESR signal response under varying DC bias for a fixed polar angle of $\theta=\pi/18$. The results are presented in Fig. 8, where a 2D color map highlights the normalized ESR signal relative to the background (baseline off-resonance) current, preserving the characteristic peak and dip behaviors. The efficiency of the ESR signal closely follows the figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ shown in Fig. 7b, demonstrating low efficiency outside the CB regime beyond the energy thresholds $\varepsilon$ and $\varepsilon+U$ for the parameter used here.
As the DC bias approaches zero, the background current diminishes, albeit at a different rate compared to $\Delta I$. This results in a non-zero ESR signal at zero bias, a phenomenon reported previously [20, 35]. This observation aligns with the earlier discussion that $\Omega_{\mathrm{FLT}}$ does not vanish at zero bias but rather at the electron-hole symmetry point, $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$. Indeed, there the signal completely disappears, as predicted by the exchange field formula, Eq. (23), and observed in Fig. 7. It is worth mentioning that if we boost the STT-driven ESR in the CB regime by forcing a larger $\hbar/\tau_{c}$, the signal at the symmetry point can become finite since it will be STT-mediated. This outcome of the model is not strictly physical, since increasing the broadening in this way is not fully consistent within a sequential tunneling picture; however, a similar effect could occur if cotunneling processes were incorporated into the model.
@@
An intriguing phenomenon observed in Fig. 8 is the shifting of the ESR resonance position. We emphasize that in these simulations, the tip-sample distance is kept constant; therefore, the couplings remain unchanged while sweeping the DC bias. This dressing of the resonance energy arises from the projection of the exchange field onto the eigenbasis axis of the spin and only vanishes at the electron-hole symmetry point or when the spin is initialized exactly perpendicular to the exchange field. In these cases, the resonance occurs at the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz.
Figure 8: Shifting of the ESR resonance position with DC bias. The ESR signal intensity, normalized to the background current, is shown as a color map, revealing peaks at positive DC biases and dips at negative DC biases. As the voltage approaches the thresholds $\varepsilon$ and $\varepsilon+U$, the ESR efficiency decreases substantially. The observed frequency shift agrees with the prediction of Eq.(LABEL:energy_dress), we have computed $\delta f/f_{0}=(f_{0}^{\prime}-f_{0})/f_{0}$, which deviates by up to approximately 8% (about 1.3 GHz) from the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz. $\Delta f=f-f_{0}$ and $\Delta I=I_{L}-I_{\mathrm{BG}}.$
This frequency shift has been consistently reported in the literature  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], with various explanations proposed. Recently, it was attributed to a change in the $g$-factor induced by the electric field of the DC bias [20]. However, experimental results [28], supported by DFT simulations, show that the net effect of the DC bias on the magnetic moment is orders of magnitude too small, ruling out $g$-factor variation as the leading cause. Ref. [23] further excludes a $g$-factor contribution: a $g$-modulation would preclude the use of Fe atoms to drive the single-spin $1/2$ Ti atom, contradicting experimental observations, a point discussed in detail by some of us in Refs. [81, 28]. This leaves the exchange field as the most plausible mechanism for the phenomenon, though minor contributions from other sources cannot be entirely ruled out.
@@
In fact, recent measurements provide supporting evidence for this mechanism [28, 30], with the latter work clearly illustrating the nonlinear region in Fig. 8. Additionally, the observable resonance frequency shift has been employed as a direct experimental method to extract the impurity parameters $\varepsilon$ and $\varepsilon+U$ [28, 30], in good agreement with DFT simulations [79]. Since the shift depicted in Fig. 8 is proportional to $\Omega_{\mathrm{FLT}}$ shown in Fig. 7a, see Eq. (31), measuring the resonance shift provides a practical way to determine the Rabi rate. This prediction could be tested using alternative Rabi-rate estimation methods, such as that presented in [25].
@@
IV.7 Vector field dependence of the Rabi rate and ESR signal: Homodyne detection
@@
The polar angle $\theta$ formed between the external magnetic field and spin polarization of the tip appears naturally in all quantities of interest and is closely related to measuring an efficient ESR signal, as we observed by comparing Figs. 5c and 6c. As such, it is important to understand the behavior of our model with respect to $\theta$. Furthermore, since the static magnetic field angle is a parameter easily adjusted in some experiments, confirming the agreement we have between model and experiment will provide validity of the driving field emerging from the modulation of the exchange magnetic field and the hopping.
@@
Figure 9: Example of the ESR signal behavior on resonance, calculating the total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, at negative (dotted line) and positive (solid line) DC bias as a function of the polar angle $\theta$ formed between the magnetic field and the tip spin polarization. The maximum ESR signal occurs at a predominantly out-of-plane angle, as a result of homodyne detection. A sizable signal for an in-plane angle arises due to the spin accumulation.
@@
As an example of the angle behavior, we have Fig. 9 showing the total ESR amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$ for the optimal DC bias values that maximize the quality factor $\Omega_{\mathrm{FLT}}T_{2}$, see Fig. 7c. $\Delta I_{s}$ denotes the symmetric part of the ESR signal, while $\Delta I_{a}$ provides the asymmetric part. Both are extracted from fitting the raw ESR signal to a symmetric and an antisymmetric Lorentzian center on the resonance frequency following Ref. [26],
$\Delta I=I_{L}-I_{\mathrm{BG}}=\Delta I_{s}\frac{1}{1+\delta^{2}}-\Delta I_{a}\frac{2\delta}{1+\delta^{2}},$ (32)
where $\delta=2(f-f_{0}^{\prime})/W=2(f-f_{0}-\delta f)/W$, with $W$ the full width at half maximum.
@@
Figure 10: Full DC and vector field dependence (Homodyne detection). a) 2D color map of the ESR total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, normalized to the background current, keeping its sign, as a function of the polar angle and DC bias, highlighting regions of high ESR efficiency. $\Delta I_{s}$ and $\Delta I_{a}$ denote symmetric and asymmetric component, see Eq. (32). At $\theta=0,\pi$ and along the electron-hole symmetry lines, no ESR signal is observed. White dashed lines indicate the cuts shown in Fig. 9. b) Figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ as a function of the polar angle $\theta$ and DC bias, showing a sinusoidal dependence within the Coulomb blockade regime proportional to $\sin\theta$, and a rapid suppression outside of it. The red dotted line marks the electron-hole symmetry line at $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$.
At $\theta=0$ or $\pi$, the ESR signal vanishes completely because the driving field and the spin are parallel. Experimentally, this is not observed, likely because the tip forms an angle with the plane containing the spin i.e. in experiments the magnetization direction of the tip is generally not exactly aligned with $z$ as in the model. This can be incorporated in the simulations by tilting the plane in which the magnetic field lies by an angle $\gamma$ relative to the spin polarization along $z$, so that the spin and tip directions do not lie in the same plane, see Appendix D.
@@
Outside $\theta=0,\pi$, the ESR amplitude at positive and negative DC bias follows the homodyne detection pattern, with a maximum out-of-plane ($\theta\sim 10$ degrees), which matches the experimental evidence [42, 22]. However, at 90 degrees, we encounter a local minimum, which goes against the conventional ESR-STM measurements, since the homodyne detection predicts zero amplitude for in-plane angles as the ESR signal is proportional to the scalar product of the spin and the magnetization direction of the tip. As we already mentioned, we identify this discrepancy as a consequence of spin accumulation, which is less pronounced in the experiments, but produces a small current contribution along $z$ axis in the model, enabling visualization of the ESR response. To suppress this feature we can lower the DC bias to move deeper into the CB regime; however, the measured current then becomes very small, Fig. 2c. In this situation, the model demands including higher-order tunneling processes (cotunneling) to produce a measurable current. It is worth noting that, since we are plotting $\Delta I_{p}$, we do not visualize the flipping of the symmetric signal happening when the angle is sweeping from 0 to $90$ degrees, going through a pure Fano profile, see Appendix D.
@@
Turning our sight to Fig. 10a, we observe the ESR amplitude dependence versus both the DC bias and $\theta$. The voltage dependence is similar to in Fig. 8 and independent of the angle, but resolving the resonance shifting since we are plotting $\Delta I_{p}$ on resonance. If we focus on the angle dependence, we see again the ESR signal vanishing in the case of parallel ($\theta=0$ degrees) and anti-parallel ($\theta=180$ degrees) orientations, with a local minimum for a perpendicular alignment ($\theta=\pm 90$ degrees) influence by the spin accumulation. As we mentioned earlier, reducing the DC bias reduces the ESR amplitude at the perpendicular alignment, even though the quality factor $\Omega_{\mathrm{FLT}}T_{2}$ maximizes at $\theta=90$ degrees, see Fig. 10b.
@@
Additionally, Fig. 10b demonstrates the sinusoidal dependence of the Rabi rate on $\theta$, as experimentally reported in [22], since $T_{2}$ is angle-independent. The electron-hole symmetry line, $eV_{\mathrm{DC}}=\varepsilon+U/2$, represented by a red dotted line in Fig. 10b, provides vanishing Rabi rate and ESR signal, consistent with Fig. 10a at $eV_{\mathrm{DC}}/|\varepsilon|=0.25$. Outside the CB regime, the quality factor drops to extremely low values, following the decrease in $T_{2}$ observed in Fig. 7a. Near zero DC bias, the ESR amplitude relative to the background exhibits a local maximum or minimum, depending on whether $\theta\in(-90^{\circ},90^{\circ})$ or falls within the range $90^{\circ},270^{\circ}]$, respectively. This ESR signal behavior aligns with the fact that the Rabi rate remains nonzero at zero bias.